### v1

# Towards a Real-World Aligned Benchmark for Unlearning in Recommender Systems

Abstract.
Modern recommender systems heavily leverage user interaction data to deliver personalized experiences. However, relying on personal data presents challenges in adhering to privacy regulations, such as the GDPR’s “right to be forgotten”. Machine unlearning (MU) aims to address these challenges by enabling the efficient removal of specific training data from models post-training, without compromising model utility or leaving residual information. However, current benchmarks for unlearning in recommender systems—most notably CURE4Rec—fail to reflect real-world operational demands. They focus narrowly on collaborative filtering, overlook tasks like session-based and next-basket recommendation, simulate unrealistically large unlearning requests, and ignore critical efficiency constraints. In this paper, we propose a set of design desiderata and research questions to guide the development of a more realistic benchmark for unlearning in recommender systems, with the goal of gathering feedback from the research community. Our benchmark proposal spans multiple recommendation tasks, includes domain-specific unlearning scenarios, and several unlearning algorithms—including ones adapted from a recent NeurIPS unlearning competition. Furthermore, we argue for an unlearning setup that reflects the sequential, time-sensitive nature of real-world deletion requests. We also present a preliminary experiment in a next-basket recommendation setting based on our proposed desiderata and find that unlearning also works for sequential recommendation models, exposed to many small unlearning requests. In this case, we observe that a modification of a custom-designed unlearning algorithm for recommender systems outperforms general unlearning algorithms significantly, and that unlearning can be executed with a latency of only several seconds.
copyright: nonejournalyear: 2025conference: FAccTRec@RecSys’25; September 22–26, 2025; Prague, Czech Republic
1. Introduction
Modern recommender systems rely heavily on user interaction data, such as clicks, ratings, and purchases to build personalized models.
The right-to-be-forgotten and machine unlearning. This personalization enhances the user experience, but at the same time makes it challenging to adhere to privacy rights such as the “right-to-be-forgotten” from the GDPR (GDPR.eu, nd) regulation in Europe and similar regulations adopted outside of Europe (California Privacy Protection Agency, nd; Dig, nd). These regulations mandate that responsible, ethical, and legally compliant AI systems give their users the right to request the timely deletion of their personal data from databases and models trained on them (Stoyanovich et al., 2022). Being able to efficiently remove data from models is also important for security-related scenarios, e.g., to quickly react to spam interactions (Buzzfeed, 2019) or unintended data leakage (Carlini et al., 2019).
On a technical level, the outlined challenges give rise to “machine unlearning” (MU), the problem of efficiently removing the influence of selected training data points from a machine learning model post training. After unlearning the data points there should be no information about these points in the model, it should ideally behave as if they had never been in the training set. Efficient machine unlearning is an active area of research (Cao and Yang, 2015; Ginart et al., 2019; Izzo et al., 2021; Schelter et al., 2021; Wang and Schelter, 2022; Wu et al., 2020a, b). The evaluation of unlearning algorithms is especially challenging since multiple aspects, such as the model utility after unlearning, the amount of information about the unlearned data still contained in the model (unlearning completeness), and the unlearning efficiency have to be considered simultaneously.
Shortcomings of existing benchmarks for unlearning in recommender systems. We argue in Section 2.1 that current unlearning benchmarks like CURE4Rec (Chen et al., 2024) do not sufficiently address the real-world demands of unlearning in recommender systems. They narrowly focus on collaborative filtering (CF) (Rendle et al., 2009; Sarwar et al., 2001; He et al., 2020; Xia et al., 2023; Ren et al., 2023), overlooking other vital tasks, such as session-based recommendation (SBR) (Hidasi et al., 2016; Jannach and Ludewig, 2017; Tan et al., 2016; Wu et al., 2019; Kang and McAuley, 2018; Li et al., 2017) and next-basket recommendation (NBR) (Rendle et al., 2010; Hu et al., 2020; Faggioli et al., 2020; Qin et al., 2021; Yu et al., 2020; Hu and He, 2019) that are common in e-commerce and frequently require unlearning capabilities. Moreover, these benchmarks simulate unlearning by deleting large chunks of data—typically 2–5% of all interactions—whereas actual systems face continuous, small-scale deletion requests, such as when individual users withdraw consent for personalization. The choice of data to unlearn is also misrepresented: rather than being random, real scenarios often evolve around specific types of content, like interactions with sensitive items (Schelter et al., 2023) or interactions from spammers (Buzzfeed, 2019). Operational efficiency is another overlooked factor; in practice, unlearning must be fast, inexpensive, and ideally executable directly on deployed models (Schelter et al., 2021). The unlearning procedures in CURE4Rec are only an order of magnitude faster than retraining, which undermines real-time responsiveness. Finally, promising general-purpose unlearning algorithms, such as those highlighted in the NeurIPS 2023 competition (Triantafillou et al., 2024), remain underexplored for recommendation-specific applications.
Towards a real-world aligned benchmark for unlearning in recommender systems. The goal of this paper is to propose the desiderata for a real-world aligned benchmark for unlearning in recommender systems, and to gather initial feedback. In Section 2.2, we formulate four extended research questions to address the shortcomings of existing benchmarks, and layout desiderata and design decisions for a new benchmark. In particular, we propose to cover NBR and SBR in addition to CF. For each recommendation task, we need different models, unlearning settings, and datasets to get insights about the unlearning performance for different data distributions. We consider various approximate unlearning algorithms for neural networks, one designed for recommendation models (Li et al., 2023b), two adapted from the NeurIPS unlearning competition (Triantafillou et al., 2024), and three designed for graph neural networks (GNNs) (Wu et al., 2023b, a; Dong et al., 2024) that are common architectures for recommender systems. As recommendation models get fully retrained in the range between once per week (Amazon Prime (Amazon Web Services, 2025)) and once per month (Tencent (Chen et al., 2025)), it is unrealistic to unlearn a large chunk of data collected over a longer period of time. In contrast to existing benchmarks, we propose to issue a number of small unlearning requests in a sequential manner instead of unlearning a single forget set all at once, as unlearning requests are usually time-sensitive. We propose to evaluate the unlearned models based on retained model utility, unlearning completeness, and unlearning efficiency.
Moreover, we present a first preliminary experiment in Section 3, designed according to our benchmark desiderata. Here, we evaluate the ability of three unlearning algorithms in a next-basket recommendation scenario, where we repeatedly unlearn the interactions with sensitive items (chosen in a domain-specific manner) from a neural model on an e-commerce dataset.
Contributions. Our contributions are as follows:
  • We discuss the shortcomings of existing benchmarks for unlearning in recommender systems (Section 2.1).
  • We propose a set of research questions and benchmark desiderata designed to address these shortcomings (Section 2.2).
  • We present preliminary experimental results for an unlearning scenario in e-commerce, designed according to our benchmark desiderata. We find that we can successfully unlearn interactions with sensitive items for sequential recommendation models, exposed to many small unlearning requests. Furthermore, we observe that general unlearning algorithms are outperformed by a custom-designed unlearning algorithm for recommender systems with certain modifications, and that unlearning can be executed with a low latency of several seconds (Section 3).
  • We make the code for our experiments available at https://github.com/pierre-lubitzsch/towards-unlearning-in-recsys, based on the repository of Li et al. (Li et al., 2023c).
2. Benchmark Proposal
We first elaborate on the shortcomings of existing unlearning benchmarks in Section 2.1 and subsequently discuss how to address these shortcomings with a new benchmark in Section 2.2.
2.1. Shortcomings of Existing Benchmarks
We argue that current benchmarks for unlearning in recommender systems, such as CURE4Rec (Chen et al., 2024), pose a valuable direction for academic research, but do not cover the characteristics and challenges of unlearning in real-world recommendation scenarios. We attribute this to the following reasons:
Recommender systems are broader than collaborative filtering. In CURE4Rec (Chen et al., 2024) only collaborative filtering is considered, however the field of recommendation is much broader. Sequential recommendation tasks use temporal information to model evolving user preferences over time, with important applications in e-commerce and retail including SBR (Hidasi et al., 2016; Jannach and Ludewig, 2017; Tan et al., 2016; Wu et al., 2019; Kang and McAuley, 2018; Li et al., 2017) and NBR (Rendle et al., 2010; Hu et al., 2020; Faggioli et al., 2020; Qin et al., 2021; Yu et al., 2020; Hu and He, 2019). A comprehensive unlearning benchmark should also cover these tasks.
Figure 1. Overview of the proposed benchmark.
Unlearning happens in small batches. In real-world applications, we expect many small unlearning requests to arrive over time, e.g., when users withdraw consent for personalization (Schelter et al., 2021) or when a security team asks to unlearn the interactions of certain spam users. It is important to handle many small repeated unlearning requests in a timely manner, ideally in a matter of seconds. However, current benchmarks only evaluate a single, unrealistically large unlearning request. In the NeurIPS unlearning competition (Triantafillou et al., 2024) and the CURE4Rec (Chen et al., 2024) benchmark, the data is unlearned at once, and the forget set size is 2% and $\geq 5\%$ respectively. Big unlearning requests like this would never appear in industry because unlearning requests are incoming sequentially over time and for realistic user counts such percentages will not be reached before the model will be retrained either way. The European e-commerce company Bol, for example, has a recommender system for >>13 million users (bol, 2025) with a total of 2.3 billion total interactions (Kersbergen et al., 2022) as input. A forget set size of 5% would correspond to unlearning the data of 650,000 users at once. It is hard to imagine why one would want to unlearn such a large number of users, and even if that were the case, the model would be retrained from scratch.
Distribution of interactions to unlearn. In some domains, the choice of interactions to unlearn for a user follows a pattern. An example is the removal of interactions with sensitive items, for instance, alcoholic items for a user suffering from addiction (Schelter et al., 2023, 2024). Another area where the interactions follow a pattern is unlearning the interactions from malicious actors such as spammers, who try to manipulate the recommender system’s behavior for a specific set of items (Buzzfeed, 2019). Current benchmarks do not evaluate domain-specific choices of the interactions to unlearn.
Efficiency needs to be a first-class citizen. Unlearning needs to be executed fast, without utility loss. In CURE4Rec (Chen et al., 2024), the runtime of unlearning is roughly one order of magnitude faster than retraining from scratch, which is not efficient enough for real-world use cases.
Wide coverage of unlearning algorithms. The only approximate unlearning method in CURE4Rec is SCIF (Li et al., 2023b), which conducts a single second-order parameter update. However, the 2023 NeurIPS unlearning competition (Triantafillou et al., 2024) contains several high scoring first-order iterative approaches to unlearning, which also need to be evaluated for recommendation tasks.
2.2. Research Questions and Benchmark Desiderata
Motivated by the discussed shortcomings, we formulate the following research questions:
  • RQ1 – How well does unlearning perform for recommendation tasks beyond collaborative filtering?
  • RQ2 – Do we need custom unlearning algorithms for recommender systems or are general algorithms for gradient-based unlearning enough?
  • RQ3 – Can we successfully handle the repeated unlearning of small batches of interactions (chosen in a domain-specific manner)?
  • RQ4 – Can we execute unlearning with sufficiently low latency to deploy it in real-world setups?
Based on these research questions, we discuss the desiderata and design of a real-world aligned benchmark for unlearning in recommender systems in this section. Figure 1 visualizes the proposed benchmark.
2.2.1. Choice of recommendation tasks and models
To address RQ1, we propose to include a wide variety of recommendation tasks and models, in addition to classical collaborative filtering.
Next-basket recommendation (NBR). In this task, we are given a history of sets of item interactions of a user and aim to predict the set of items they will choose next. A common application of this is predicting the next shopping basket of a user in e-commerce or retail based on their past purchases. We plan to include two kNN-based models (TIFUKNN (Hu et al., 2020), UP-CF@r (Faggioli et al., 2020)), the attention-based model Sets2Sets (Hu and He, 2019), CLEA (Qin et al., 2021), which applies constrastive learning and the GNN-based model DNNTSP (Yu et al., 2020).
Session-based recommendation (SBR). This task evolves around predicting the next item that a user will interact with, based on a sequence of past interactions. Major applications are to predict the next product to show to a user in e-commerce or to predict the next song that a user will listen to on a music streaming platform. We plan to include the kNN-based model S-kNN (Jannach and Ludewig, 2017), two RNN-based models (GRU4Rec (Tan et al., 2016), NARM (Li et al., 2017)), the attention-based model SASRec (Kang and McAuley, 2018) and the GNN-based model SR-GNN (Wu et al., 2019).
Collaborative filtering (CF). A classical task in recommender systems, already covered in CURE4Rec, where the goal is to predict the rating that a user would give to a particular unseen item, based on their past preferences. We include classical approaches such as Matrix Factorization (Rendle et al., 2009) and item-based kNN (Sarwar et al., 2001), as well as three graph-based models (LightGCN (He et al., 2020), SimRec (Xia et al., 2023), DCCF (Ren et al., 2023)).
2.2.2. Choice of datasets
We leverage a selection of widely used, publicly available datasets for our benchmark, as detailed in Table 1. The three NBR datasets are grocery shopping or retail related. Two of the three datasets for SBR are click datasets from e-commerce, while the third one is built from social media posts about songs listened to. For CF, we choose three common datasets containing ratings from users about items, extending the range of domains covered in CURE4Rec.
Table 1. Choice of datasets across various recommendation tasks for the proposed benchmark.
Dataset Domain Task #Users/Sessions #Items #Interactions Available at
TaFeng (Das, 2025) Grocery shopping NBR 32,266 23,812 817,741 Kaggle
Dunnhumby (dunnhumby, 2025) Retail shopping NBR 2,500 92,339 2,595,732 dunnhumby.com
Instacart (Le, 2025) Grocery shopping NBR 19,435 13,897 3,346,083 GitHub
RSC15 (Hidasi et al., 2016) E-commerce interactions SBR 9,249,729 52,739 33,003,944 Kaggle
DIGI (dig, [n. d.]) E-commerce interactions SBR 310,325 156,813 1,235,381 Codalab
NOWP (Zangerle et al., 2014) Music on social media SBR 738,200 3,734,862 1,291,251,677 Zenodo
Goodreads (pypiahmad, 2025) Book ratings CF 876,145 2,360,650 228,648,342 Kaggle
MovieLens (Harper and Konstan, 2015) Movie ratings CF 138,493 26,744 20,000,263 Kaggle
Amazon Reviews (Hou et al., 2024) Product reviews CF 54,510,000 48,190,000 571,540,000 GitHub
2.2.3. Choice of unlearning algorithms
To address RQ2, we propose to include recommendation-specific as well as general unlearning algorithms in the benchmark. Unlearning algorithms not used during the preliminary experiments in this paper can be found in Appendix B.
Custom unlearning algorithms for recommender systems. We do not include general exact unlearning algorithms like RecEraser (Chen et al., 2022) or UltraRE (Li et al., 2023a), because their running time depends on the training set size, which is by design not efficient enough to handle a lot of queries with small forget sets. We focus on approximate unlearning algorithms with a time complexity linear in the forget set size. We propose to include the following algorithms:
  • SCIF, a method to directly remove the influence of data in the forget set utilizing influence functions. The influence of the data in the forget set is calculated similarly to a Newton step, which additionally includes retain data to prevent catastrophic forgetting via
    $\tilde{\theta}_{z\to\bar{z}}=\tilde{\theta}-H_{\tilde{\theta}}^{-1}\nabla_{\tilde{\theta}}(\frac{1}{2+bs}(-\ell(z,\theta)+\ell(\bar{z},\theta)+\sum_{i=1}^{bs}\ell(z_{i},\theta)))$
    where $z_{i}$ are retain samples, $z$ is an unlearning sample, $\bar{z}$ is the modified unlearning sample, $H$ is the Hessian, $bs$ is the number of retain samples used for the update, and $\tilde{\theta}$ is a subset of the parameters $\theta$ of the model (Li et al., 2023b).
General unlearning algorithms for gradient-based machine learning. We additionally propose to evaluate the following unlearning algorithms from the NeurIPS’23 unlearning challenge (Triantafillou et al., 2024), which have not specifically been designed for recommendation models. We select these particular approaches, as they scored the highest in the final ranking of the competition:
  • Fanchuan: This method first operates on the forget set and learns to make the distribution for the output of the model on a unlearning sample uniform by minimizing the KL-Divergence between the output of the model on the unlearning samples and a uniform pseudo-label. In the next stage, the method varies between two modes for a certain number of epochs: in the first mode, we have a contrastive loss maximizing the distance between representations of samples in the forget set and samples in the retain set. In the other mode, we simply train the model on a subset of the retain data using the same procedure as training the model in a standard manner.
  • Kookmin: This method takes the model and calculates the gradients of the loss w.r.t. the parameters of the model on the forget set and a subset of the retain set. This is done to identify parameters to reset, having a similar gradient in the forget and retain set. With this information, the method chooses the top $p\,|\theta|$ parameters $\hat{\theta}$ having the smallest absolute gradient difference for gradients from the forget and retain set: $\hat{\theta}=(\theta_{\bm{i}})_{\bm{i}\in\mathcal{J}}$ with $p\in(0,1)$ being a hyperparameter. Here $\mathcal{J}=\operatorname*{arg\,min}_{\mathcal{X}\subseteq\text{ind}(\theta),|\mathcal{X}|=\lfloor p|\theta|\rfloor}\sum_{\bm{i}\in\mathcal{X}}|(\nabla_{\theta}\ell(\theta,D_{f}))_{\bm{i}}-(\nabla_{\theta}\ell(\theta,B_{r}))_{\bm{i}}|$ is the index set choosing the subset of these parameters where $\text{ind}(t)$ references the index set of a tensor $t$ and $B_{r}$ is a batch of the retain set of size $O(|D_{f}|)$. These parameters will be re-initialized as if they were newly created, for example a parameter $w$ in a linear layer is re-initialized using uniform initialization (He et al., 2015) and other parameters $w$ are initialized by sampling from a centered normal distribution with low standard deviation: $w\sim\mathcal{N}(0,0.02^{2})$. After that, the model is trained on a subset of the retain set, where the parameters that were not re-initialized have a smaller learning rate by one order of magnitude.
2.2.4. Evaluation metrics
We discuss the dimensions and metrics to evaluate the performance of unlearning algorithms.
Recommendation utility. We propose to leverage common ranking metrics such as Recall, Personalized hit ratio (PHR) and Normalized Discounted Cumulative Gain (nDCG).
Unlearning completeness. Next to efficient unlearning and performance of the unlearned model, it is important to measure how much the model unlearned about the forget set. For exact unlearning methods, such evaluation is unnecessary because all information about the forget set is removed by definition. Measuring unlearning completeness for approximate methods is not trivial, because most approximate unlearning algorithms have no known theoretical guarantees (see the framework of $(\varepsilon,\delta)$-unlearning outlined in Appendix A). Without theoretical guarantees, there are several methods to measure unlearning empirically. Standard ones use the retrained model and calculate the distance in weight space or the distance in output distribution of the retrained and unlearned model. Another approach to measuring unlearning completeness uses a hypothesis testing interpretation of unlearning with a fixed $\delta$ to get an estimated $\varepsilon$ for $(\varepsilon,\delta)$-unlearning (Triantafillou et al., 2024).
Sensitive item unlearning. A task specific evaluation method is sensitive item unlearning (Schelter et al., 2023). In this setting we want to remove items from a sensitive category for specific users, for example meat for users changing to a vegetarian diet, or alcohol for users suffering from addiction. For the evaluation, we count how many users still get sensitive items recommended after their unlearning request, in comparison to the recommendations of a retrained model. To evaluate recommendation utility we remove sensitive items from the test data of affected users.
Removal of poisonous data. Another evaluation method involves removing poisonous data from malicious actors in session-based recommendation systems. These actors inject fake clicks on random items and items they want to promote, corrupting the training data and causing models to recommend irrelevant items. Unlearning success can be measured using standard performance metrics, as the goal is eliminating the negative impact of these fake interactions.
We propose to use both empirical setups to measure unlearning completeness and to construct corresponding experiments, which repeatedly unlearn small portions of the data, to account for RQ3.
Unlearning efficiency. To account for RQ4, we propose to measure the time it takes to execute unlearning requests, as well as the required memory and hardware (e.g., if accelerator hardware is needed). Another important aspect is whether offline precomputation (independent of the choice of the forget set) is possible. In that case, we propose to report the precomputation time and the execution time of the unlearning algorithm separately.
3. Preliminary Results
In our preliminary experiment, we focus on the task of NBR and construct an experiment according to our desiderata from Section 2.2.
Table 2. Results for sensitive item unlearning on Sets2sets (Instacart dataset, averaged over five seeds). Kookmin and Fanchuan effectively remove sensitive items, with Fanchuan being fastest. SCIF performs badly in utility and completeness. However, when including norm clipping of the update step for SCIF, its performance drastically improves (SCIF clip). Bold/underlined entries show best/second-best performance across approximate algorithms for % of requests.
Category Requests Algorithm Recall nDCG PHR Time per Sensitive items KL(R——U)
@10 @20 @10 @20 @10 @20 request (s) @10 @20
alcohol 25% Fanchuan 0.2071 0.3026 0.2088 0.2379 0.6596 0.7725 6.31 0.00% 0.00% 3.7439
alcohol 50% Fanchuan 0.2008 0.2909 0.2025 0.2294 0.6474 0.7632 6.31 0.00% 0.00% 4.2762
alcohol 75% Fanchuan 0.2010 0.2891 0.2040 0.2296 0.6467 0.7615 6.31 0.00% 0.07% 4.5580
alcohol 100% Fanchuan 0.2022 0.2878 0.2061 0.2302 0.6493 0.7601 6.31 0.04% 0.39% 4.6788
alcohol 25% Kookmin 0.2091 0.3005 0.2254 0.2511 0.6836 0.7811 46.35 0.00% 0.00% 0.9800
alcohol 50% Kookmin 0.2078 0.2984 0.2233 0.2488 0.6816 0.7792 46.35 0.00% 0.00% 1.0526
alcohol 75% Kookmin 0.2060 0.2961 0.2213 0.2468 0.6787 0.7769 46.35 0.00% 0.05% 1.1253
alcohol 100% Kookmin 0.2072 0.2973 0.2212 0.2468 0.6807 0.7776 46.35 0.00% 0.06% 1.1269
alcohol 25% SCIF 0.1097 0.2080 0.0852 0.1244 0.4502 0.6626 55.40 25.09% 30.50% 17.3190
alcohol 50% SCIF 0.1111 0.2092 0.0837 0.1229 0.4555 0.6646 55.40 17.86% 28.44% 19.8368
alcohol 75% SCIF 0.1161 0.2106 0.0921 0.1289 0.4654 0.6631 55.40 17.66% 27.47% 33.4127
alcohol 100% SCIF 0.1184 0.2131 0.0944 0.1313 0.4736 0.6686 55.40 17.75% 27.90% 38.3761
alcohol 25% SCIF clip 0.2162 0.3170 0.2210 0.2518 0.7152 0.8282 95.01 0.00% 0.00% 0.6006
alcohol 50% SCIF clip 0.2135 0.3144 0.2197 0.2509 0.7069 0.8226 95.01 0.00% 0.00% 0.7062
alcohol 75% SCIF clip 0.2128 0.3143 0.2206 0.2520 0.7055 0.8235 95.01 0.00% 0.00% 0.8098
alcohol 100% SCIF clip 0.2133 0.3141 0.2220 0.2529 0.7052 0.8222 95.01 0.00% 0.00% 0.8810
alcohol 25% Retrain 0.2220 0.3172 0.2295 0.2567 0.6968 0.7922 2977.40 0.00% 0.00% 0.0000
alcohol 50% Retrain 0.2219 0.3172 0.2297 0.2570 0.6963 0.7925 3059.40 0.00% 0.00% 0.0000
alcohol 75% Retrain 0.2216 0.3173 0.2299 0.2572 0.6971 0.7937 2939.20 0.00% 0.00% 0.0000
alcohol 100% Retrain 0.2225 0.3178 0.2306 0.2577 0.6982 0.7937 3025.60 0.00% 0.00% 0.0000
meat 25% Fanchuan 0.2061 0.3006 0.2091 0.2374 0.6598 0.7716 6.00 0.36% 2.88% 3.2123
meat 50% Fanchuan 0.2045 0.2984 0.2056 0.2340 0.6552 0.7695 6.00 0.31% 4.80% 3.4694
meat 75% Fanchuan 0.2012 0.2923 0.2029 0.2302 0.6489 0.7644 6.00 0.00% 0.52% 3.8409
meat 100% Fanchuan 0.2009 0.2886 0.2038 0.2293 0.6474 0.7613 6.00 0.09% 1.49% 3.8590
meat 25% Kookmin 0.2111 0.3018 0.2276 0.2528 0.6877 0.7825 46.15 0.12% 1.31% 0.8741
meat 50% Kookmin 0.2078 0.2980 0.2239 0.2491 0.6829 0.7793 46.15 0.06% 1.33% 0.9681
meat 75% Kookmin 0.2075 0.2977 0.2229 0.2484 0.6818 0.7790 46.15 0.24% 2.27% 1.0024
meat 100% Kookmin 0.2067 0.2965 0.2223 0.2476 0.6809 0.7777 46.15 0.33% 3.16% 1.0248
meat 25% SCIF 0.0965 0.1641 0.0819 0.1072 0.3680 0.5051 38.50 49.96% 65.80% 25.7528
meat 50% SCIF 0.0956 0.1634 0.0772 0.1030 0.3669 0.5040 38.50 46.92% 61.09% 31.9693
meat 75% SCIF 0.0958 0.1635 0.0772 0.1030 0.3677 0.5039 38.50 46.74% 61.28% 34.0601
meat 100% SCIF 0.0599 0.1087 0.0475 0.0665 0.2422 0.3561 38.50 47.10% 72.02% 35.4680
meat 25% SCIF clip 0.2176 0.3170 0.2220 0.2521 0.7188 0.8286 95.36 0.00% 0.00% 0.6509
meat 50% SCIF clip 0.2147 0.3166 0.2189 0.2505 0.7099 0.8249 95.36 0.00% 0.00% 0.7082
meat 75% SCIF clip 0.2142 0.3149 0.2189 0.2503 0.7078 0.8235 95.36 0.00% 0.42% 0.7154
meat 100% SCIF clip 0.2130 0.3142 0.2182 0.2497 0.7053 0.8221 95.36 0.00% 0.16% 0.7664
meat 25% Retrain 0.2221 0.3183 0.2296 0.2573 0.6976 0.7935 3076.00 0.00% 0.60% 0.0000
meat 50% Retrain 0.2227 0.3195 0.2303 0.2582 0.6980 0.7948 3051.80 0.00% 0.48% 0.0000
meat 75% Retrain 0.2222 0.3185 0.2301 0.2578 0.6973 0.7942 3067.20 0.28% 3.18% 0.0000
meat 100% Retrain 0.2222 0.3182 0.2301 0.2576 0.6975 0.7939 2601.80 0.00% 0.82% 0.0000
Experimental setup. We choose Sets2sets (Hu and He, 2019) as a model, which was among the best performing models in Li et al. (Li et al., 2023c), and leverage the Instacart dataset, which contains transactions from grocery shopping. For unlearning, we evaluate the approximate neural unlearning techniques SCIF (Li et al., 2023b) and the two top approaches from the 2023 NeurIPS unlearning competition (Triantafillou et al., 2024).
We evaluate the performance of the algorithms for sensitive item unlearning, with alcohol and meat as sensitive item categories. We train the models on our dataset with five different seeds. To construct a forget set for a sensitive category $c$, we randomly choose $0.1\%$ of the user-item interactions from items in baskets and sample only items from category $c$. This results in $\sim 1000$ users per forget set for alcohol and $\sim 650$ users per forget set for meat. For each seed we sample a different forget set. Let $D_{f}^{c,s}$ contain all forget requests sampled for category $c$ with seed $s$. We omit indices $c,s$ in further notation for readability. To unlearn $D_{f}$, we first fix an order of the users submitting a forget request: $(u_{1},\dots,u_{\ell})$. Let $D$ be the initial training set, $D_{f}^{i}$ be unlearning request $i$ from $u_{i}$, $D_{r}^{i}=D\setminus\bigcup_{j\leq i}D_{f}^{j}$ be the retain set for request $i$, and $\theta_{i}$ be the weights of the model after unlearning $D_{f}^{1},\dots,D_{f}^{i}$. We sequentially execute the unlearning algorithm on input $(\theta_{i-1},D_{r}^{i},D_{f}^{i})$ for $i\in\{1,\dots,\ell\}$. We save $\theta_{i}$ at every quarter of the total forget set size $\ell$, so that we can evaluate the results after different amounts of unlearning requests. For each (category, seed) combination, we unlearn one model per unlearning algorithm. We retrain models from scratch on $D_{r}^{i}$ for four checkpoints $i\in C:=\{\lceil\frac{\ell}{4}\rceil,\lceil\frac{2\ell}{4}\rceil,\lceil\frac{3\ell}{4}\rceil,\lceil\frac{4\ell}{4}\rceil\}$ and all (category, seed) combinations to compare the retrained models $\theta_{i}^{r}$ to the unlearned models $\theta_{i}$ for $i\in C$. Each user $u_{j}$ has one input sample $z$ consisting of their basket history. We use a temporal-order leave-one-out split to form a train/val/test split. When executing the unlearning algorithm on input $(\theta_{i-1},D_{r}^{i},D_{f}^{i})$, we unlearn all sensitive items from the baskets of $u_{i}$.
Besides common utility metrics, we report the fraction of users that still receive recommendations for sensitive items after unlearning and the KL divergence between retrained and unlearned models for users who submitted unlearning requests, as a measure of unlearning completeness. To quantify efficiency, we report the average time to unlearn the interactions of a single user. We conduct each experiment with a single NVIDIA A100 80GB PCIe GPU.
Necessary adaptations for unlearning algorithms. The NBR setting requires us to adapt the unlearning algorithms. Fanchuan originally uses the whole retain set for the repair round. This would make the algorithm too inefficient in our case, and we limit the retain samples used per retain round to $10\,|D_{f}^{i}|$ and only use 16 instead of 256 retain samples during the contrastive stage. Kookmin conducts a reset of 30% of the model parameters in the setup for the NeurIPS competition; however, in our setting of sequential unlearning of small unlearning sets, it makes sense to lower this hyperparameter. We chose a value of $1\%$ of the parameters of all layers because 30% made the model collapse due to shorter retain rounds, and $0.1\%$ rendered the updates negligible. We have 5 retain rounds and train on $32\,|D_{f}^{i}|$ samples each round, instead of training 5 rounds on the full retain set. We adapt this to scale with the forget set for the sake of efficiency. Further studies on the best value of the reset percentage are needed. SCIF by default only updates the parameters for the user embeddings in CF models. However, in NBR, the information about a user is usually implicit within the basket history, and the chosen models do not contain user embeddings. Therefore, we make SCIF update the item embeddings, output linear layer, and a linear layer after attention for Sets2sets.
[…]
Preliminary results for this can be seen in rows with Algorithm SCIF clip. Due to time constraints these results only stem from using one seed instead of averaging over five seeds, and $c=1$ is chosen heuristically. Further experiments need to be done to find the best value of $c$ and to confirm the results for this seed. SCIF clip outperforms the other algorithms in model utility and unlearning completeness metrics. However, SCIF clip is an order of magnitude slower than Fanchuan and Fanchuan might perform better than SCIF clip given more time.
The general unlearning algorithms Fanchuan and Kookmin perform well. They achieve utility levels comparable to the retrained models, with a similarly effective reduction in sensitive item predictions. Kookmin performs slightly better in most cases. Furthermore, we observe that the average time to unlearn the interactions of a single user is in the range of six seconds for Fanchuan. From the KL divergence over the forget set, we can see that the predictions of Kookmin are always closer in distribution to the retrained model than the predictions from Fanchuan. In summary, this experiment provides first evidence that our research questions from Section 2.2 can be positively answered with a new benchmark. The experiment indicates that unlearning also works for NBR models (RQ1) with many small unlearning requests (RQ3). We observe a case where a custom-designed unlearning algorithm for recommender systems outperforms general unlearning algorithms (RQ2). This only holds when applying update step norm clipping for the custom-designed unlearning algorithm. Without this modification, the performance drops significantly. Furthermore, unlearning algorithms can be executed with a low latency of several seconds only, which demonstrates the practical potential for deploying unlearning in real-world production systems (RQ4).
4. Next Steps
Our plan for the future is to evolve our benchmark proposal (based on community feedback) into a comprehensive, fully specified, and openly available public benchmark, conduct all proposed experiments, and report on the results. We hope that our benchmark helps lay the foundation to deploy efficient unlearning for users in real-world systems.
Acknowledgements.
This research was (partially) supported by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union’s Horizon Europe program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.
Appendix A Background on Machine Unlearning
Machine unlearning is the process of removing the influence of specific data from a machine learning model as if that data had never been seen during training. For the formal definition of unlearning consider the following: Let $D$ be the original training dataset, $\mathcal{A}$ be a training algorithm, $U$ be an unlearning algorithm, $D_{f}\subseteq D$ be the forget set to unlearn, and $D_{r}=D\setminus D_{f}$ be the retain set, e.g. the data retained after unlearning. $U$ is an exact unlearning algorithm iff for all $T\subseteq\mathcal{H}$:
$\mathbb{P}(\mathcal{A}(D_{r})\in T)=\mathbb{P}(U(\mathcal{A}(D),D_{r},D_{f})\in T)$
where $\mathcal{H}$ is the hypothesis space of models. The intuition behind the formula is that the probability of receiving a model $\mathcal{A}(D_{r})$ after retraining from scratch on the retain set $D_{r}$ should be the same as the probability of receiving this model from the unlearning procedure $U$ starting at the model $\mathcal{A}(D)$ unlearning $D_{f}$. This implies the same distribution of weights and model outputs. For theoretical guarantees for approximate unlearning there is the notion of $(\varepsilon,\delta)$-unlearning (Triantafillou et al., 2024), where $U$ is a $(\varepsilon,\delta)$-unlearning algorithm iff for all $T\subseteq\mathcal{H}$:
$\mathbb{P}(\mathcal{A}(D_{r})\in T)\leq\exp(\varepsilon)\,\mathbb{P}(U(\mathcal{A}(D),D_{r},D_{f})\in T)+\delta$
and
$\mathbb{P}(U(\mathcal{A}(D),D_{r},D_{f})\in T)\leq\exp(\varepsilon)\,\mathbb{P}(\mathcal{A}(D_{r})\in T)+\delta$
with $\varepsilon\in[0,\infty),\delta\in[0,1)$. Here, $\varepsilon$ is a multiplicative error margin, and $\delta$ is an additive error margin, where lower values of these parameters denote stronger unlearning bounds. Note that $(0,0)$-unlearning is equivalent to exact unlearning.
Appendix B Additional unlearning algorithms
  • GIF, which works similarly to SCIF, but the parameters considered for the parameter update are chosen differently: The node embedding of node $u$ in the graph is contained in the parameter update if there exists a node $v$ such that $v$ is in the forget set and the distance from $u$ to $v$ is smaller than a hyperparameter $d$ (Wu et al., 2023b).
  • CEU, a framework to unlearn edges from a GNN. To get theoretical guarantees for linear models they first add a linear noise term $b^{T}\theta$ with $b\sim\mathcal{N}(0,\sigma^{2}I)$ to the training loss and fine-tune aiming to hide the real gradient residual. After that, they unlearn using an influence function on all parameters (Wu et al., 2023a).
  • IDEA, which analyses the original loss over the whole training set and the pruned loss which excludes nodes and edges we want to unlearn. It then conducts a second-order update step based on the difference between the original and pruned objective loss (Dong et al., 2024).
  • Caboose, an exact unlearning algorithm for neighborhood-based recommendation on sparse data, which uses efficient indexing and incremental view maintenance to update nearest neighbor sets with sub-second latency (Schelter et al., 2023, 2024).